OpenStudy (anonymous):

1. What is the equation for slope-intercept form? A. y = mx + b B. bx + y = m C. y = bx + m

OpenStudy (anonymous):

a

OpenStudy (anonymous):

A

hartnn (hartnn):

slope - intercept ---> y=mx+b

hartnn (hartnn):

m=slope, y-intercept = b

OpenStudy (anonymous):

thank you i thought it was either that or c i couldn't quite remember which letter was where

OpenStudy (anonymous):

. Write the equation of the line through (2, -1) with a slope of 4 in point-slope form. A. y + 1 = 4(x - 2) B. y - 1 = 4( x + 2) C. y + 2 = 4( x -1) D. y + 4 = -1(x- 2) 8. What is the formula for standard form of an equation? A. Bx + Ay = C B. Cx + Ay = B C. Ax + By = C

OpenStudy (anonymous):

ill write these down after for extra help

OpenStudy (anonymous):

B for the first one and C for the last one

OpenStudy (anonymous):

. Write an equation of the line that passes through (-5, 2) and (6, 2). A. y = -5x + 6 B. y = 2x + 2 C. y = 0x + 2 or y= 2 D. y = -5x + 2 6. True or False: The following formula is in point-slope form. A. true B. false

OpenStudy (anonymous):

C and im not sure if its true of false

OpenStudy (anonymous):

Its C

OpenStudy (anonymous):

formula for # 6 is 

OpenStudy (anonymous):

Thats point slope form

OpenStudy (anonymous):

A. y - 1 = -2/3(x - 4) B. 2x + 3y = 11 C. -2/3 x + 4y = 11 D. 3y = -2x + 11

OpenStudy (anonymous):

b looks like the only thats in standard form but i could be wrong

OpenStudy (anonymous):

whats the question?

OpenStudy (anonymous):

A. true B. false

OpenStudy (anonymous):

the question is in the attachments it a picture.

OpenStudy (anonymous):

true

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

A. Solve for b B. Plug in the slope C. Calculate slope D. Write the final equation

OpenStudy (anonymous):

what is the last one asking for?

OpenStudy (anonymous):

Given two points, what is the first step to write an equation in slope-intercept form?

OpenStudy (anonymous):

then it would be B

OpenStudy (anonymous):

solve for b that is